# ExpGest: Expressive Speaker Generation Using Diffusion Model and Hybrid Audio-Text Guidance

### v1

@@ Abstract
Existing gesture generation methods primarily focus on upper body gestures based on audio features, neglecting speech content, emotion, and locomotion. These limitations result in stiff, mechanical gestures that fail to convey the true meaning of audio content. We introduce ExpGest, a novel framework leveraging synchronized text and audio information to generate expressive full-body gestures. Unlike AdaIN or one-hot encoding methods, we design a noise emotion classifier for optimizing adversarial direction noise, avoiding melody distortion and guiding results towards specified emotions. Moreover, aligning semantic and gestures in the latent space provides better generalization capabilities. ExpGest, a diffusion model-based gesture generation framework, is the first attempt to offer mixed generation modes, including audio-driven gestures and text-shaped motion. Experiments show that our framework effectively learns from combined text-driven motion and audio-induced gesture datasets, and preliminary results demonstrate that ExpGest achieves more expressive, natural, and controllable global motion in speakers compared to state-of-the-art models.
Index Terms:
Gesture generation, multimodel learning, emotion guide, locmotion control
@@ I Introduction
In fields like virtual agents, movies, and human-computer interaction, virtual speakers convey information through co-speech gestures related to audio melody and content, and non-spontaneous movements. The emerging field of audio gesture generation has gained attention, with early research using rule-based methods [1], while data-driven techniques [2] improved diversity using statistical models. Deep models, such as VAE [3], flow models [4], and diffusion-based models, generate gestures directly from raw audio data. Methods combining audio melody and semantics [5] have advanced significantly. However, DiffStyleGesture [6] and Emog [7], which use emotion as guidance, perform poorly on the BEAT dataset [8].
@@
In non-spontaneous motion generation, early methods [9] considered deterministic signal-to-motion mapping using neural networks. The complexity and randomness of motion led to deeper research on generative models. TEMOS [10] pioneered action-to-action and text-to-action tasks. Recently, KIT-ML [11] and HumanML3D [12] datasets promoted text-guided motion generation. T2M-GPT [13] uses autoregressive models, improving results but with jitter issues. Diffusion model-based methods (e.g., MDM [14], MLD [15]) enhance motion data continuity through denoisers. Despite advancements, bottlenecks persist. No work has coherently integrated both motion categories, and different motion representations lead to dataset inconsistency, keeping the field in the single modality generation stage.
@@ II-C Semantic Alignment in Latent Space
where, $z_{s}$ and $z_{g}$ are the latent space representations of a matching transcription-gesture pair. $sim$ is similarity score between two latent codes, $K$ is a set containing one positive sample transcription and a group of negative sample gestures, and $\tau$ is the temperature parameter used to adjust the sensitivity of the function. Finally, we freeze the trained semantic alignment module and deploy only the transcription encoder into the GDM, ensuring that the final generated results accurately capture the semantic content.
@@ II-D Noise-based Emotion Guided Classifier
$\hat{x_{t}}=x_{t}+\alpha\cdot\nabla_{xt}\mathcal{L}(Emo\_Classifier(x_{t}),y),$ (6)
@@ III-A Dataset and Evaluation Metrics
TABLE I: Quantitative Assessment Outcomes. Boldface denotes the top-performing method for each evaluation criterion. All experiments are conducted and appraised on the BEAT dataset.
Method $FGD(Raw)$ $FGD(Feature)$ $EA$ $SA$ Ground Truth 0.0 0.0 0.97 0.91 CAMN [8] 52.4 263.9 - - Trimodal [23] 47.6 212.7 - - DiffStyleGesture [6] 33.7 133.9 0.60 0.11 ExpGest 11.7 76.6 0.91 0.61 ExpGest w/ hybrid-gudied 25.4 129.3 - -
Method $EA$ $EC$ $EA_{hands}$ $EC_{hands}$ Ground Truth 0.97 - 0.95 - DiffStyleGesture [6] 0.60 0.27 0.49 0.19 DiffStyleGesture w/ EG 0.83 0.69 0.70 0.63 ExpGest 0.91 0.83 0.81 0.70
Method Human-likeness Gesture-appropriateness Emotion-compatibility Global-coherence Ground Truth $4.61\pm 0.17$ $4.72\pm 0.20$ $3.62\pm 0.27$ DiffStyleGesture [6] $3.46\pm 0.16$ $3.61\pm 0.11$ $1.22\pm 0.19$ ExpGest (audio-only) $4.17\pm 0.13$ $4.20\pm 0.18$ $3.88\pm 0.21$ ExpGest (audio-action) $3.76\pm 0.08$ $3.83\pm 0.11$ $2.63\pm 0.09$ $3.61\pm 0.17$ ExpGest (audio-text) $3.72\pm 0.10$ $3.91\pm 0.15$ $2.27\pm 0.06$ $3.47\pm 0.16$
@@ III-D User Study and Results
Results. Figure 4 (top) demonstrates a comparison of ExpGest with prior state-of-the-art methods under single audio guidance. We uniformly selected serene recordings from Wayne and Solomon, two speakers in the BEAT dataset, as test audio. The generation results of DiffStyleGesture, which relies solely on melody features, are relatively static with little variation. In contrast, ExpGest, benefiting from semantic alignment in the latent space, reveals highly expressive gesture results. Furthermore, our proposed noise emotion guide can naturally and smoothly perform emotional transitions, thereby enhancing the expressiveness of the generated results for reference). Figure 4 (bottom) showcases the superior performance of ExpGest under mixed control. By integrating audio with action, ExpGest can generate a variety of motion sequences, such as “dragging left leg, running forward”. Simultaneously, using existing large language models to edit the action text description for a speech (“he looks for a place to sit while talking”), combined with the audio itself, a vivid speaking character can be generated.
III-E Ablation Studies
@@ IV Conclusions
In this paper, we propose ExpGest, a novel framework using the diffusion generation model to create motion speakers from mixed audio-text guidance. Training on heterogeneous gesture-audio and text-motion pairs allows effective capture of human motion nuances. Our approach generates natural-looking speaker gesture sequences, laying the groundwork for large-scale motion speaker generation with potential applications in virtual agents, movies, and human-computer interaction.

### v2

@@ Abstract
Existing gesture generation methods primarily focus on upper body gestures based on audio features, neglecting speech content, emotion, and locomotion. These limitations result in stiff, mechanical gestures that fail to convey the true meaning of audio content. We introduce ExpGest, a novel framework leveraging synchronized text and audio information to generate expressive full-body gestures. Unlike AdaIN or one-hot encoding methods, we design a noise emotion classifier for optimizing adversarial direction noise, avoiding melody distortion and guiding results towards specified emotions. Moreover, aligning semantic and gestures in the latent space provides better generalization capabilities. ExpGest, a diffusion model-based gesture generation framework, is the first attempt to offer mixed generation modes, including audio-driven gestures and text-shaped motion. Experiments show that our framework effectively learns from combined text-driven motion and audio-induced gesture datasets, and preliminary results demonstrate that ExpGest achieves more expressive, natural, and controllable global motion in speakers compared to state-of-the-art models. Our code, model, and demo are are available at https://github.com/cyk990422/ExpGest/.
@@ I Introduction
In fields like virtual agents, movies, and human-computer interaction, virtual speakers convey information through co-speech gestures related to audio melody and content, and non-spontaneous movements. The emerging field of audio gesture generation has gained attention, with early research using rule-based methods [16], while data-driven techniques [11] improved diversity using statistical models. Deep models, such as VAE [15], flow models [27], and diffusion-based models [26, 6], generate gestures directly from raw audio data. Methods combining audio melody and semantics [1] have advanced significantly. However, DiffStyleGesture [25] and Emog [28], which use emotion as guidance, perform poorly on the BEAT dataset [18].
@@
In non-spontaneous motion generation, early methods [9] considered deterministic signal-to-motion mapping using neural networks. The complexity and randomness of motion led to deeper research on generative models. TEMOS [22] pioneered action-to-action and text-to-action tasks. Recently, KIT-ML [23] and HumanML3D [10] datasets promoted text-guided motion generation. T2M-GPT [31] uses autoregressive models, improving results but with jitter issues. Diffusion model-based methods [24, 4, 12, 13, 7] enhance motion data continuity through denoisers. Despite advancements, bottlenecks persist. No work has coherently integrated both motion categories, and different motion representations lead to dataset inconsistency, keeping the field in the single modality generation stage.
@@ II-C Semantic Alignment in Latent Space
where, $z_{s}$ and $z_{g}$ are the latent space representations of a matching transcription-gesture pair. simsim is similarity score between two latent codes, $K$ is a set containing one positive sample transcription and a group of negative sample gestures, and $\tau$ is the temperature parameter used to adjust the sensitivity of the function. Finally, we freeze the trained semantic alignment module and deploy only the transcription encoder into the GDM, ensuring that the final generated results accurately capture the semantic content.
@@ II-D Noise-based Emotion Guided Classifier
$\hat{x}_{t}=x_{t}+\alpha\cdot\nabla_{x_{t}}\mathcal{L}(\text{Emo\_Classifier}(x_{t}),y)$ (6)
@@ III-A Dataset and Evaluation Metrics
Method $FGD(Raw)$ $FGD(Feature)$ EAEA SASA Ground Truth 0.0 0.0 0.97 0.91 CAMN [18] 52.4 263.9 - - Trimodal [29] 47.6 212.7 - - DiffStyleGesture [25] 33.7 133.9 0.60 0.11 ExpGest 11.7 76.6 0.91 0.61 ExpGest w/ hybrid-gudied 25.4 129.3 - -
TABLE I: Quantitative Assessment Outcomes. Boldface denotes the top-performing method for each evaluation criterion. All experiments are conducted and appraised on the BEAT dataset. Trimodal [29] and CAMN [18] outcomes are reproduced utilizing the official code, while DiffStyleGesture [25] results are retrained on the BEAT [18] dataset employing the official training approach. A dash (“-”) signifies the unavailability of a particular metric.
@@ III-B Comparison to State-of-the-art Methods
Method EAEA ECEC $EA_{hands}$ $EC_{hands}$ Ground Truth 0.97 - 0.95 - DiffStyleGesture [25] 0.60 0.27 0.49 0.19 DiffStyleGesture w/ EG 0.83 0.69 0.70 0.63 ExpGest 0.91 0.83 0.81 0.70
Method Human-likeness Gesture-appropriateness Emotion-compatibility Global-coherence Ground Truth $4.61\pm 0.17$ $4.72\pm 0.20$ $3.62\pm 0.27$ - DiffStyleGesture [25] $3.46\pm 0.16$ $3.61\pm 0.11$ $1.22\pm 0.19$ - ExpGest (audio-only) $4.17\pm 0.13$ $4.20\pm 0.18$ $3.88\pm 0.21$ - ExpGest (audio-action) $3.76\pm 0.08$ $3.83\pm 0.11$ $2.63\pm 0.09$ $3.61\pm 0.17$ ExpGest (audio-text) $3.72\pm 0.10$ $3.91\pm 0.15$ $2.27\pm 0.06$ $3.47\pm 0.16$
@@ III-D User Study and Results
Results. Figure 4 (top) demonstrates a comparison of ExpGest with prior state-of-the-art methods under single audio guidance. We uniformly selected serene recordings from Wayne and Solomon, two speakers in the BEAT dataset, as test audio. The generation results of DiffStyleGesture, which relies solely on melody features, are relatively static with little variation. In contrast, ExpGest, benefiting from semantic alignment in the latent space, reveals highly expressive gesture results. Furthermore, our proposed noise emotion guide can naturally and smoothly perform emotional transitions, thereby enhancing the expressiveness of the generated results (see Spp video for reference). Figure 4 (bottom) showcases the superior performance of ExpGest under mixed control. By integrating audio with action, ExpGest can generate a variety of motion sequences, such as “dragging left leg, running forward”. Simultaneously, using existing large language models to edit the action text description for a speech (“he looks for a place to sit while talking”), combined with the audio itself, a vivid speaking character can be generated.
III-E Ablation Studies
@@ IV Conclusions
In this paper, we propose ExpGest, a novel framework using the diffusion generation model to create motion speakers from mixed audio-text guidance. Training on heterogeneous gesture-audio and text-motion pairs allows effective capture of human motion nuances. Our approach generates natural-looking speaker gesture sequences, laying the groundwork for large-scale motion speaker generation with potential applications in virtual agents, movies, and human-computer interaction. In the future, we will incorporate existing action capture methods [5, 17] to generate more diverse continuous natural data [30], so as to achieve more realistic training effects.